Number of real roots of the equation esin x –e-sin x -4 = 0 is

Grade:12

2 Answers

Gurwinder Kaur
13 years ago

esin x.esin x-4esin x-1=0

esin x=(4+2\sqrt{5}. \,  )/2 , (4-2\sqrt{5}. \,  )/2

=2+\sqrt{5}. \,  ,2-\sqrt{5}. \,

approve if u like....

mounika gajawada
13 years ago

ans: zero

on solvin tha quadratic in e^(sinx) we obtain values 2+√5 and 2-√5 => sinx = ln(2+√5)/ln(2-√5)

ln(2+√5) > 1 and ln(2-√5)- is nt defined

hence der is no real root  for given equation.....